# Approximate Polymorphisms of Predicates

### v1

@@ 4 Main theorem
The proof also relies on a two-sided version of It Ain’t Over Till It’s Over, which immediately follows from the one-sided version, Theorem 3.2, by applying it to both $f$ and $1-f$.
@@ 4.1 Counting lemma
The proof of the case $\chi_{\epsilon}\equiv*$ will require the following lemma, whose proof is adapted from [Mos12a].
@@
The remainder of the argument is reminiscent of the corresponding argument in [CFM+22]. As in the proof of the counting lemma for NAND (Lemma 3.5), we will sample $\mu$ using a two-step process. Let $u_{\ast}$ be obtained from $u$ by setting the $j_{0}$’th coordinate to *. First, we sample a restriction $\rho\in(\{0,1\}^{m}\cup\{u_{\ast}\})^{n}$ by sampling each coordinate independently according to the following law, where $q=\min(\mu(u),\mu(u^{\oplus j_{0}}))$:
@@ 4.2 Rounding
Let $P\subseteq\{0,1\}^{m}$ be a non-empty coordinate. A coordinate $j\in[m]$ is flexible if there exists $w_{(j)}\in P$ such that $w_{(j)}^{\oplus j}\in P$, and inflexible otherwise.
@@
If $j$ is inflexible then we need to use the fact that every extension of $\chi_{\epsilon}$ belongs to $P$. Since $j$ is inflexible, this implies that $\chi_{\epsilon}(j)\neq*$. We consider two subcases, according to the value of $\chi_{\epsilon}(j)$. If $\chi_{\epsilon}(j)=0$ then $\operatorname*{\mathbb{E}}[f_{j}|_{J\leftarrow x^{(j)}}]\leq\epsilon$. If $\chi_{\epsilon}(j)=1$ then $\operatorname*{\mathbb{E}}[f_{j}|_{J\leftarrow x^{(j)}}]\geq 1-\epsilon$. Since $j$ is inflexible, $\nu|_{j}=\mu|_{j}$, and so $f_{j}|_{J\leftarrow x^{(j)},\rho}$ is a constant, which equals 00 with probability at most $\epsilon$. In both cases, the summand is bounded by $\epsilon$. ∎
@@ 4.3 Proof without short affine relations
For every $j\in I\setminus F$, apply Lemma 4.11 with $\mu:=\mu|_{j}$, $\eta:=\eta^{\prime}$, $\zeta:=1/(2m)$, and the $q$ in the definition of $\nu$ (Section 4.2) to obtain $M_{j}$. The lemma implies that if $|S_{j}\setminus J|\geq M_{j}$ then for every $x^{(j)}\in\{0,1\}^{J}$ we have $\eta^{\prime}<\operatorname*{\mathbb{E}}[\chi_{S_{j},b_{j}}|_{J\leftarrow x^{(j)},\rho|_{j}}]<1-\eta^{\prime}$ with probability $1-1/(2m)$ over the choice of $\rho$.
@@
Apply Lemma 4.6 (Rounding lemma) with the predicate $P|_{I}$, the distribution $\mu|_{I}$, $\epsilon:=\eta^{\prime}$, and $\zeta:=1/(2m)$ to obtain $\delta^{\prime\prime}$ such that the following holds. If $(f^{\prime}_{1},\dots,f^{\prime}_{m})$ is a $(\mu,\min(\mu)^{|J|}\delta^{\prime\prime})$-approximate polymorphism of $P|_{I}$ then $(g_{j}^{\rho,\eta^{\prime}})_{j\in I}$ is a generalized polymorphism of $P|_{I}$ w.p. $1-1/(2m)$ over the choice of $\rho$.
@@
Lemma 4.6 implies that the first property holds with probability $1-1/(2m)$. We go on to the second property. Let $j\in I\setminus F$. If $S_{j}\subseteq J$ then $g^{\rho,\eta}_{j}=f^{\prime}_{j}$ always. Otherwise, $|S_{j}\setminus J|\geq M_{j}$, and so Lemma 4.11 implies that $g^{\rho,\eta}_{j}=f^{\prime}_{j}$ with probability at least $1-1/(2m)$ (crucially, the property $\chi_{j}^{\rho,\eta}=*$ doesn’t depend on the input $x^{(j)}$ to $J$). In total, the first two properties hold with probability at least $1/2$.
@@ 4.4 Concluding the proof
Let $Z$ be the set of coordinates in $P$ that are always 00. Let $I$ consist of a choice of one coordinate from each set of equivalent non-constant coordinates. For each $j\in I$, let $I_{j}$ be the coordinates equivalent to $j$; possibly $I_{j}=\{j\}$. The sets $Z,(I_{j})_{j\in I}$ thus partition $[m]$.
@@ 5 Linearity testing
In this section we prove Theorem 1.6, which extends the BLR test to arbitrary distributions. The proof uses ideas from [DFH25], which extended a related test to a specific family of distributions.
See 1.6
@@
The general case is handled by reduction to the uniform distribution, using an agreement theorem from [DFH25]. The basic idea is to sample $x\sim\mu$ in two steps. In the first step, we sample $x\sim\nu$ with probability $1-q$ (for appropriate $\nu,q$), and leave it unsampled with probability $q$. In the second case, if $x$ wasn’t sampled, we sample it according to $\pi$.
@@
Our argument will require two agreement theorems: one from [DFH25], and a folklore result about mixing Markov chains which we prove in Section 5.2. The reduction itself is described in Section 5.3, where we also provide a more detailed outline.
@@ 5.1 Uniform distribution
In this section we prove Theorem 1.6 in the case of the uniform distribution. The proof follows the classical argument of [BCH+96].
@@
where $B=(-1)^{b}$ and $x^{(1)},\dots,x^{(m)}$ are sampled as follows: the first $m-1$ are sampled uniformly and independently, and $x^{(m)}_{i}=Bx^{(1)}_{i}\cdots x^{(m-1)}_{i}$.
@@ 5.3 Arbitrary distributions
    Using the agreement theorem of [DFH25] we show that for each $\alpha\in P_{m,b}^{n}$ there exists a consensus set $A(\alpha)$ such that typically $A(S,\alpha|_{S^{c}})=A(\alpha)\cap S$.
@@ 5.3.1 Step 1
The proof will require an agreement theorem essentially proved in [DFH25]. The theorem concerns the distribution $\mu_{q,r}^{n}$, where $10<q<r<$ 1. This is the distribution on triples $(S_{1},S_{2},T)$ defined as follows:
@@
Let $10<r<q<$ 1. Suppose that for every $S\subseteq[n]$ we have a function $\phi_{S}\colon S\to\Sigma$, where $\Sigma$ is some finite alphabet. If
@@
The slice version of this result is [DFH25, Theorem 3.1]. This version can be proved using the “going to infinity” argument which is used to prove [DFH25, Theorem 5.4]. ∎
@@
We take $r=(1-c)q$. We will show that
@@ 6 Intersecting families
In this section we prove Theorem 1.9, which improves on results of Friedgut and Regev [FR18] (their paper, in turn, improves on [DFR08, DF09]).
@@
The proof relies on the machinery of Friedgut and Regev, which reduces it to the following statement.
Theorem 6.1 (Fractional improved Friedgut–Regev).
Fix $p\in(0,1/2)$. Let $P_{\mathsf{NAND}}=\{(0,0),(1,0),(0,1)\}$, and let $\mu$ be the following distribution: $\mu(0,0)=1-2p$, $\mu(1,0)=\mu(0,1)=p$. For every $\epsilon>0$ there exist $\delta=1/\exp\Theta(1/\epsilon^{C})$ and $M=O(1/\epsilon^{C})$ (for some global constant $C$) such that the following holds.
@@
Then there exist $g_{1},g_{2}\colon\{0,1\}^{n}\to\{0,1\}$, computed by a decision tree of depth at most $M$, such that $(g_{1},g_{2})$ is a generalized polymorphism of $P_{\mathsf{NAND}}$, and moreover for $j\in\{1,2\}$,
@@
Apply Lemma 6.4 with $\epsilon:=\epsilon/2$ to obtain $d=\Theta(\log(1/\epsilon))$ and $\tau,\gamma=\Theta(1/\epsilon^{C})$.
We apply Theorem 6.3 with $\epsilon:=\epsilon/2$ and the parameters $d,\tau$ to $f_{1},f_{2}$ to obtain a decision tree $T$ of depth $M=O(d/\epsilon\tau)=\tilde{O}(1/\epsilon^{2C})$. We define the functions $g_{1},g_{2}\colon\{0,1\}^{n}\to\{0,1\}$ as follows. For each leaf $\rho\in T$,
$g_{j}|_{\rho}=\begin{cases}0&\text{if $f_{j}|_{\rho}$ is not $(d,\tau)$-regular or $\operatorname*{\mathbb{E}}_{\mu|_{j}}[f_{j}|_{\rho}]\leq\epsilon/2$},\\ 1&\text{otherwise}.\end{cases}$
If we sample $\rho$ according to $\mu|_{j}$ then according to Jones’ regularity lemma, $f_{j}|_{\rho}$ is $(d,\tau)$-regular with probability at least $21-\epsilon/$ 2. This shows that $\operatorname*{\mathbb{E}}_{\mu|_{j}}[(1-g_{j})f_{j}]\leq\epsilon/2+\epsilon/2=\epsilon$.
It remains to show that $(g_{1},g_{2})$ is a generalized polymorphism of $P_{\mathsf{NAND}}$ for small enough $\delta$. If this is not the case, then there exists a partial assignment $\rho\in T$ such that $(g_{1}|_{\rho|_{1}},g_{2}|_{\rho|_{2}})$ is not a generalized polymorphism of $P_{\mathsf{NAND}}$, which implies that $f_{1}|_{\rho|_{1}},f_{2}|_{\rho|_{2}}$ are both regular and have expectation at least $\epsilon/2$. Lemma 6.4 thus implies that $\operatorname*{\mathbb{E}}_{\mu}[f_{1}|_{\rho|_{1}}f_{2}|_{\rho|_{2}}]\geq\gamma$, which is ruled out by defining $\delta=\min(\mu)^{M}\gamma/2=1/\exp\tilde{\Theta}(1/\epsilon^{2C})$. ∎
@@ 7 General alphabets
Let $\rho$ be a random restriction obtained by sampling each coordinate $i\in[n]$ independently according to the following law: with probability $1-q$, draw a sample from $\mu$, and otherwise, draw *.
@@ 8 Regularity lemmas
We derive all of them from the following two versions.
Theorem 8.1 (Jones’ regularity lemma (decision tree version)).
For every alphabet $\Sigma$, every $m\in\mathbb{N}$, every full support distributions $\mu_{1},\dots,\mu_{m}$, and every $\epsilon,\tau>0$, $d\in\mathbb{N}$ the following holds for $M=O(md/\epsilon\tau)$.
For all functions $f_{1},\dots,f_{m}\colon\Sigma^{n}\to[0,1]$ and every decision tree $T_{0}$ there exists a decision tree $T$ extending $T_{0}$ with additional depth at most $M$ such that for all $j$,
$\Pr_{\rho\sim T}[f_{j}|_{\rho}\text{ is $(d,\tau)$-regular with respect to $\mu_{j}$}]\geq 1-\epsilon,$
where $\rho$ is sampled by following $T$, sampling each variable encountered according to $\mu_{j}$.
Theorem 8.2 (Jones’ regularity lemma (junta version)).
@@
For all functions $f_{1},\dots,f_{m}\colon\Sigma^{n}\to[0,1]$ and set $J_{0}\subseteq[n]$ there exists set $J\supseteq J_{0}$ of size at most $\mathcal{M}(|J|)$ such that for all $j$,
@@
Theorems 3.3 and 6.3 follow from Theorem 8.1, and Theorems 4.2 and 7.3 follow from Theorem 8.2.
The proofs proceed via a different notion of regularity, using noisy influences.
@@
Theorem 8.4 (Jones’ regularity lemma for noisy influences (decision tree version)).
For every alphabet $\Sigma$, every $m\in\mathbb{N}$, every full support distributions $\mu_{1},\dots,\mu_{m}$, and every $\epsilon,\tau>0$, $d\in\mathbb{N}$ the following holds for $M=m/\epsilon\tau\cdot\rho/(1-\rho)$.
For all functions $f_{1},\dots,f_{m}\colon\Sigma^{n}\to[0,1]$ and every decision tree $T_{0}$ there exists a decision tree $T$ extending $T_{0}$ with an additional depth at most $M$ such that for all $j$,
$\Pr_{\rho\sim T}[f_{j}|_{\rho}\text{ is $(\rho,\tau)$-noisy-regular with respect to $\mu_{j}$}]\geq 1-\epsilon,$
where $\rho$ is sampled by following $T$, sampling each variable encountered according to $\mu_{j}$.
The proof of Theorem 8.4 uses a potential argument. For a decision tree $T$, we define
@@
Proof of Theorem 8.4.
We construct a sequence of trees $T_{t}$ as follows. The starting point is $T_{0}$. If $T_{t}$ doesn’t satisfy the conclusion of the lemma for $f_{j}$, then Lemma 8.5 implies that we can add one level to $T_{t}$, forming a tree $T_{t+1}$ satisfying
$\Phi^{\rho}_{f_{j},\mu_{j}}(T_{t+1})\geq\Phi^{\rho}_{f_{j},\mu_{j}}(T_{t})+\frac{1-\rho}{\rho}\epsilon\tau.$
For $k\neq j$, the lemma gives the guarantee
$\Phi^{\rho}_{f_{k},\mu_{k}}(T_{t+1})\geq\Phi^{\rho}_{f_{k},\mu_{k}}(T_{t}),$
and so
$\Phi(T_{t+1})\geq\Phi(T_{t})+\frac{1-\rho}{\rho}\epsilon\tau.$
Since $\operatorname{Stab}_{\rho}[f]\leq\|f\|^{2}$ for every function $f$, it is easy to check that $0\leq\Phi(T)\leq m$ for all $T$. In particular, the process above must terminate after at most $m/\epsilon\tau\cdot\rho/(1-\rho)$ steps. Taking the last tree to be our $T$ completes the proof. ∎
Proof of Theorem 8.1.
The proof of Theorem 8.2 similarly follows from a version for noisy influences.
Theorem 8.6 (Jones’ regularity lemma for noisy influences (junta version)).
$\Pr_{x\sim\mu_{j}}[f_{j}|_{J\leftarrow x}\text{ is $(\rho,\tau)$-noisy-regularwith respect to $\mu_{j}$ for all $j$}]\geq 1-\epsilon.$
The proof is a modification of the proof of Theorem 8.4.
@@
As in the proof of Theorem 8.4, the process terminates within $1/\epsilon\tau\cdot\rho/(1-\rho)$ steps. Since $|J_{t+1}|\leq|J_{t}|+|\Sigma|^{|J_{t}|}$, the final size of the resulting set $J$ can be bounded independently of $n$. ∎
The proof makes it clear that the dependence of $\mathcal{M}$ on the parameters is much worse in Theorem 4.2 than in Theorem 3.3. For this reason, it would be interesting to find a proof of Theorem 1.5 which uses the decision tree version rather than the junta version.

### v2

@@ 4 Main theorem
The proof also relies on a two-sided version of It Ain’t Over Till It’s Over, which immediately follows from the one-sided version, Theorem 3.2, by applying it to both $f$ and 1f1-f.
@@ 4.1 Counting lemma
The proof of the case $\chi_{\epsilon}\equiv*$ will require the following lemma, whose proof is adapted from [Mossel2012].
@@
The remainder of the argument is reminiscent of the corresponding argument in [CFMMS2022]. As in the proof of the counting lemma for NAND (Lemma 3.5), we will sample $\mu$ using a two-step process. Let $u_{\ast}$ be obtained from $u$ by setting the $j_{0}$’th coordinate to *. First, we sample a restriction $\rho\in(\{0,1\}^{m}\cup\{u_{\ast}\})^{n}$ by sampling each coordinate independently according to the following law, where $q=\min(\mu(u),\mu(u^{\oplus j_{0}}))$:
@@ 4.2 Rounding
Let $P\subseteq\{0,1\}^{m}$ be a non-empty predicate. A coordinate $j\in[m]$ is flexible if there exists $w_{(j)}\in P$ such that $w_{(j)}^{\oplus j}\in P$, and inflexible otherwise.
@@
If $j$ is inflexible then we need to use the fact that every extension of $\chi_{\epsilon}$ belongs to $P$. Since $j$ is inflexible, this implies that $\chi_{\epsilon}(j)\neq*$. We consider two subcases, according to the value of $\chi_{\epsilon}(j)$. If $\chi_{\epsilon}(j)=0$ then $\operatorname*{\mathbb{E}}[f_{j}|_{J\leftarrow x^{(j)}}]\leq\epsilon$. If $\chi_{\epsilon}(j)=1$ then $\operatorname*{\mathbb{E}}[f_{j}|_{J\leftarrow x^{(j)}}]\geq 1-\epsilon$. Since $j$ is inflexible, $\nu|_{j}=\mu|_{j}$, and so $f_{j}|_{J\leftarrow x^{(j)},\rho}$ is a constant, which equals 0 with probability at most $\epsilon$. In both cases, the summand is bounded by $\epsilon$. ∎
@@ 4.3 Proof without short affine relations
For every $j\in I\setminus F$, apply Lemma 4.11 with $\mu:=\mu|_{j}$, $\eta:=\eta^{\prime}$, $\zeta:=1/(2m)$, and the $q$ in the definition of $\nu$ (Section 4.2) to obtain $M_{j}$. The lemma implies that if $|S_{j}\setminus J|\geq M_{j}$ then for every $x^{(j)}\in\{0,1\}^{J}$ we have $\eta^{\prime}<\operatorname*{\mathbb{E}}[\chi_{S_{j},b_{j}}|_{J\leftarrow x^{(j)},\rho|_{j}}]<1-\eta^{\prime}$ with probability 11/(2m)1-1/(2m) over the choice of $\rho$.
@@
Apply Lemma 4.6 (Rounding lemma) with the predicate $P|_{I}$, the distribution $\mu|_{I}$, $\epsilon:=\eta^{\prime}$, and $\zeta:=1/(2m)$ to obtain $\delta^{\prime\prime}$ such that the following holds. If $(f^{\prime}_{1},\dots,f^{\prime}_{m})$ is a $(\mu,\min(\mu)^{|J|}\delta^{\prime\prime})$-approximate polymorphism of $P|_{I}$ then $(g_{j}^{\rho,\eta^{\prime}})_{j\in I}$ is a generalized polymorphism of $P|_{I}$ w.p. 11/(2m)1-1/(2m) over the choice of $\rho$.
@@
Lemma 4.6 implies that the first property holds with probability 11/(2m)1-1/(2m). We go on to the second property. Let $j\in I\setminus F$. If $S_{j}\subseteq J$ then $g^{\rho,\eta}_{j}=f^{\prime}_{j}$ always. Otherwise, $|S_{j}\setminus J|\geq M_{j}$, and so Lemma 4.11 implies that $g^{\rho,\eta}_{j}=f^{\prime}_{j}$ with probability at least 11/(2m)1-1/(2m) (crucially, the property $\chi_{j}^{\rho,\eta}=*$ doesn’t depend on the input $x^{(j)}$ to $J$). In total, the first two properties hold with probability at least $1/2$.
@@ 4.4 Concluding the proof
Let $Z$ be the set of coordinates in $P$ that are always 0. Let $I$ consist of a choice of one coordinate from each set of equivalent non-constant coordinates. For each $j\in I$, let $I_{j}$ be the coordinates equivalent to $j$; possibly $I_{j}=\{j\}$. The sets $Z,(I_{j})_{j\in I}$ thus partition $[m]$.
@@ 5 Linearity testing
In this section we prove Theorem 1.7, which extends the BLR test to arbitrary distributions. The proof uses ideas from [DFH2025], which extended a related test to a specific family of distributions.
See 1.7
@@
The general case is handled by reduction to the uniform distribution, using an agreement theorem from [DFH2025]. The basic idea is to sample $x\sim\mu$ in two steps. In the first step, we sample $x\sim\nu$ with probability 1q1-q (for appropriate $\nu,q$), and leave it unsampled with probability $q$. In the second case, if $x$ wasn’t sampled, we sample it according to $\pi$.
@@
Our argument will require two agreement theorems: one from [DFH2025], and a folklore result about mixing Markov chains which we prove in Section 5.2. The reduction itself is described in Section 5.3, where we also provide a more detailed outline.
@@ 5.1 Uniform distribution
In this section we prove Theorem 1.7 in the case of the uniform distribution. The proof follows the classical argument of [BCHKS96].
@@
where $B=(-1)^{b}$ and $x^{(1)},\dots,x^{(m)}$ are sampled as follows: the first m1m-1 are sampled uniformly and independently, and $x^{(m)}_{i}=Bx^{(1)}_{i}\cdots x^{(m-1)}_{i}$.
@@ 5.3 Arbitrary distributions
    Using the agreement theorem of [DFH2025] we show that for each $\alpha\in P_{m,b}^{n}$ there exists a consensus set $A(\alpha)$ such that typically $A(S,\alpha|_{S^{c}})=A(\alpha)\cap S$.
@@ 5.3.1 Step 1
The proof will require an agreement theorem essentially proved in [DFH2025]. The theorem concerns the distribution $\mu_{q,r}^{n}$, where $0<q<r<1$. This is the distribution on triples $(S_{1},S_{2},T)$ defined as follows:
@@
Let $0<r<q<1$. Suppose that for every $S\subseteq[n]$ we have a function $\phi_{S}\colon S\to\Sigma$, where $\Sigma$ is some finite alphabet. If
@@
The slice version of this result is [DFH2025, Theorem 3.1]. This version can be proved using the “going to infinity” argument which is used to prove [DFH2025, Theorem 5.4]. ∎
@@
We take r=(1c)qr=(1-c)q. We will show that
@@ 6 Intersecting families
In this section we provide an alternative proof Theorem 1.9, originally due to Friedgut and Regev [FR18] (who improved earlier results of Dinur, Friedgut, and Regev [DFR08, DR09]).
@@
Instead of working directly with functions on $\binom{[n]}{pn}$, Friedgut and Regev first reduce Theorem 1.9 to the following statement.
Theorem 6.1 (Fractional Friedgut–Regev).
Fix $p\in(0,1/2)$. Let $P_{\mathsf{NAND}}=\{(0,0),(1,0),(0,1)\}$, and let $\mu$ be the following distribution: $\mu(0,0)=1-2p$, $\mu(1,0)=\mu(0,1)=p$. For every $\epsilon>0$ there exist $\delta,M>0$ such that the following holds.
@@
Then there exist functions $g_{1},g_{2}\colon\{0,1\}^{n}\to\{0,1\}$, depending on at most $M$ coordinates, such that $(g_{1},g_{2})$ is a generalized polymorphism of $P_{\mathsf{NAND}}$, and moreover for $j\in\{1,2\}$,
@@
Apply Lemma 6.4 with $\epsilon:=\epsilon/2$ to obtain $d,\tau,\gamma$.
We apply Theorem 6.3 with $\epsilon:=\epsilon/2$ and the parameters $d,\tau$ to $f_{1},f_{2}$ to obtain a set $J$ of size $M$. We define the functions $g_{1},g_{2}\colon\{0,1\}^{n}\to\{0,1\}$ as follows. For each $x\in\{0,1\}^{J}$,
$g_{j}|_{J\leftarrow x}=\begin{cases}0&\text{if $f_{j}|_{J\leftarrow x}$ is not $(d,\tau)$-regular or $\operatorname*{\mathbb{E}}_{\mu|_{j}}[f_{j}|_{J\leftarrow x}]\leq\epsilon/2$},\\ 1&\text{otherwise}.\end{cases}$
If we sample $x$ according to $\mu|_{j}$ then according to Jones’ regularity lemma, $f_{j}|_{J\leftarrow x}$ is $(d,\tau)$-regular with probability at least $1-\epsilon/2$. This shows that $\operatorname*{\mathbb{E}}_{\mu|_{j}}[(1-g_{j})f_{j}]\leq\epsilon/2+\epsilon/2=\epsilon$.
It remains to show that $(g_{1},g_{2})$ is a generalized polymorphism of $P_{\mathsf{NAND}}$ for small enough $\delta$. If this is not the case, then there exists an assignment $(x^{(1)},x^{(2)})\in P^{J}$ such that $(g_{1}|_{J\leftarrow x^{(1)}},g_{2}|_{J\leftarrow x^{(2)}})$ is not a generalized polymorphism of $P_{\mathsf{NAND}}$, which implies that $f_{1}|_{J\leftarrow x^{(1)}},f_{2}|_{J\leftarrow x^{(2)}}$ are both regular and have expectation at least $\epsilon/2$. Lemma 6.4 thus implies that $\operatorname*{\mathbb{E}}_{\mu}[f_{1}|_{J\leftarrow x^{(1)}}f_{2}|_{J\leftarrow x^{(2)}}]\geq\gamma$, which is ruled out by defining $\delta=\min(\mu)^{M}\gamma/2$. ∎
@@ 7 General alphabets
Let $\rho$ be a random restriction obtained by sampling each coordinate $i\in[n]$ independently according to the following law: with probability 1q1-q, draw a sample from $\mu$, and otherwise, draw *.
@@ 8 Regularity lemmas
We derive all of them from the following general version.
Theorem 8.1 (Jones’ regularity lemma).
@@
For all functions $f_{1},\dots,f_{m}\colon\Sigma^{n}\to[0,1]$ and set $J_{0}\subseteq[n]$ there exists a set $J\supseteq J_{0}$ of size at most $\mathcal{M}(|J|)$ such that for all $j$,
@@
The proof proceeds via a different notion of regularity, which uses noisy influences.
@@
Theorem 8.3 (Jones’ regularity lemma for noisy influences).
[…]
The proof of Theorem 8.3 uses a potential argument, and employs decision trees as intermediate representations. For a decision tree $T$, we define
@@
Since $\operatorname{Stab}_{\rho}[f]\leq\|f\|^{2}$ for every function $f$, it is easy to check that $0\leq\Phi(T)\leq m$ for all $T$. In particular, the process above must terminate after at most $m/\epsilon\tau\cdot\rho/(1-\rho)$ steps. Since $|J_{t+1}|\leq|J_{t}|+|\Sigma|^{|J_{t}|}$, the final size of the resulting set $J$ can be bounded independently of $n$. ∎
Proof of Theorem 8.1.
The proof makes it clear that the dependence of $\mathcal{M}$ on the parameters is quite bad (of tower type).